OpenStudy (anonymous):

Given that: (1+x)^5 (1+ax)^6 ≡ 1 + bx + 10x^2 + ..... + a^6x^11, find the values of a,b∈Z I'm not entirely sure how to go about doing this. Please help? :)

OpenStudy (asnaseer):

1. use the binomial expansion to expand each bracket up to the the terms involving \(x^2\) 2. then multiply these two keeping only the terms up to \(x^2\) (i.e. ignore any terms of higher powers of x) 3. then compare the result with \(1 + bx + 10x^2\) and you should get some equations involving a and b that can be solved.

OpenStudy (anonymous):

Thanks I'll do the expansion and see if I'm on the right track.

OpenStudy (asnaseer):

great! let me know if you get stuck.

OpenStudy (anonymous):

when I do the expansion, I get = 1+(30ax^2)+(150*a^2*x^4) Am I on the right track? thanks

OpenStudy (asnaseer):

what is this the expansion for?

OpenStudy (anonymous):

for (1+x)^5 (1+ax)^6

OpenStudy (asnaseer):

That does not look correct. I suggest you do this step-by-step. first show what you believe will be the expansion for (1+x)^5 then the expansion for (1+ax)^6 then we can work on the product

OpenStudy (anonymous):

Yeah, it seems I've done too many steps at once and got confused. I'll work on them separately. :)

OpenStudy (anonymous):

I get: (1+x)^5 = 1 + 5x + 10x^2 (1+ax)^6 = 1 + 6ax + 15ax^2

OpenStudy (asnaseer):

I /think/ the 3rd term in the 2nd expansion should have \(a^2\) in it

OpenStudy (asnaseer):

i.e.:\[(1+ax)^6 = 1 + 6ax + 15a^2x^2\]

OpenStudy (anonymous):

Ah yse, forgot the brackets.

OpenStudy (asnaseer):

think of "ax" as one quantity

OpenStudy (asnaseer):

ok, good, now you need to calculate this beast:\[(1 + 5x + 10x^2)(1 + 6ax + 15a^2x^2)\]

OpenStudy (anonymous):

Should I do it term by term?

OpenStudy (anonymous):

first multiply the 1 by each of the three terms, then the 5x ... etc.?

OpenStudy (asnaseer):

yes - I would break it up into three parts as follows:\[1(1 + 6ax + 15a^2x^2)+5x(1 + 6ax + 15a^2x^2)+10x^2(1 + 6ax + 15a^2x^2)\]

OpenStudy (asnaseer):

and remember to throw away all powers of x greater than 2

OpenStudy (anonymous):

ah so I'm allowed to just disregard them? That's pretty handy. :D

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

great thanks :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

throwing away the powers above x^2, I get the product to be: \[1 + 6ax + 15a ^{2}x ^{2}+5x+30ax ^{2}+10x ^{2}\]

OpenStudy (asnaseer):

good - now gather like terms and you almost there.

OpenStudy (anonymous):

\[1 + (6a+5)x + (15a ^{2}+30a+10)x ^{2}\] so b = 6a+5 and the other bracket equals to 10?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

you should get two solutions for 'a'

OpenStudy (anonymous):

I get a = (b-5)/6 and a = -2

OpenStudy (anonymous):

oh, I plug -2 into the first a equation to get b?

OpenStudy (asnaseer):

a=-2 is correct but you missed another value for a. once you know the values for a, plug them into b=6a+5 to get the values for b.

OpenStudy (anonymous):

I got that a(a+2) = 0 so a = 0 or -2, but since a belongs to Z then it's just -2? Or is 0 a valid value too then?

OpenStudy (asnaseer):

I thought Z included all negative, zero, and positive numbers?

OpenStudy (asnaseer):

so a=0 is a solution as well

OpenStudy (anonymous):

Ah yes, probably. Thanks very much for the help! To think this isn't even the toughest question on my assigment. ... -.- :)

OpenStudy (asnaseer):

yw :) and good luck with the rest!

OpenStudy (anonymous):

Cheers!

OpenStudy (asnaseer):

np my friend! :)